OpenStudy (anonymous):

Which of the following is equivalent to log 35 8 rounded to three decimal places? 1.710 0.714 0.585 –0.641

OpenStudy (anonymous):

not clear is it \(\log(358)\) or \(\log_{35}(8)\)

OpenStudy (anonymous):

second one

OpenStudy (anonymous):

then with a calculator compute \[\frac{\ln(8)}{\ln(35)}\] to get your answer by the change of base formula

OpenStudy (anonymous):

i got it thanks